OpenStudy (anonymous):

an airplane is sighted at the same time by two ground observers who are 2 miles apart and in line with the airplane. they report the angles of elevation as 10 degrees and 22 degrees. how high is the airplane?

OpenStudy (anonymous):

Let A, B, C be the position of the first and the second observer and the point directly beneath the airplane. Let H be the height of the air plane. Then we have tan 10deg=h/AC tan 22 deg=h/BC AC-BC=2 This can be solved for the three unknowns h, AC and BC. I find h=0.625 miles. Does that make sense or do you need a further explanation?

OpenStudy (anonymous):

further explanation

OpenStudy (anonymous):

 that was my drawing

OpenStudy (mathmale):

Caroline, Would you mind saying what it is that you'd still like to learn in regard to Kirby's solution?